# Uncertainty Quantification and Confidence Calibration in Large Language Models: A Survey

### v1

@@ 2.2. Uncertainty and Confidence in LLMs
As uncertainty and confidence are often intertwined, many approaches used in uncertainty quantification have their counterpart in confidence estimation. For example, for black-box methods, Lin et al. (2023) computes a similarity matrix of sampled responses and derives confidence estimates for each generation via its degree or distance derived from the graph Laplacian, before using these scores to compute uncertainty. Zhang et al. (2024a) extends such black-box methods to longer generations. For logit-based methods, Malinin and Gales (2021) normalize Eq. 1 with the length of $\mathbf{s}$. Further improvements include replacing the logit-sum or mean with weighted sum, by attention values on the important tokens (Lin et al., 2024) or by importance inferred from natural language inference (NLI) models (Duan et al., 2024). Such variants of sequence likelihood could then be fed for (entropy-style) uncertainty computation (Kuhn et al., 2023; Lin et al., 2024).
@@
Another popular approach is asking the LM itself whether a particular free-form generation is correct (Kadavath et al., 2022). However, this formulation also poses a restriction on the confidence estimation method, as it is essentially a scalar logit. Thus, many extensions focus on applying calibration methods from classification to calibrate such self-evaluation. The few exceptions include Ren et al. (2023b), which generalizes Kadavath et al. (2022) and converts samples from free-form generation into a multiple-choice question (with generations being the options) and adding a ”None of the above” option to elicit the confidence. Similarly, Shrivastava et al. (2025) performs the explicit relative comparison of pairs of answers and then aggregates the preferences into an absolute confidence level.
@@
Since we typically care about the LM’s confidence in the “semantic space” due to semantic invariance, instead of manipulating logits, a popular approach is to perform additional training for confidence estimation. This could be done on the base LM (either full LM (Zhao et al., 2023a; Jiang et al., 2021; Kapoor et al., 2024) or partial (Liu et al., 2024a)) with a different loss, or using a separate model on the internal or external representations from the base LM (Jagannatha and Yu, 2020; Mielke et al., 2022; Ren et al., 2023a; Azaria and Mitchell, 2023; Ulmer et al., 2024). On the other end of the spectrum, without any training, prompting could be used to elicit verbalized confidence values (Xiong et al., 2024a; Tian et al., 2023), or to recalibrate LLM confidence for a particular distribution (Li et al., 2024b) via in-context learning. Finally, one could combine multiple confidence estimation methods and enjoy the benefit of ensembling (Gao et al., 2024).
@@
Just like the evaluation for uncertainty quantification (more in Section 4), the choice of correctness function has a profound impact on the conclusion of the experiments, especially for free-form generation tasks. Popular choices include using (potentially larger) LLM as judges (Liu et al., 2024a; Lin et al., 2023; Tian et al., 2023), human annotations (Ren et al., 2023b; Mielke et al., 2022), or lexical similarities such as ROUGE (Kuhn et al., 2023; Zhao et al., 2023a). Recently, Liu et al. (2025) proposes to evaluate free-form generation confidence measures with selected multiple-choice datasets as an efficient complement. For longer generations, Huang et al. (2024a) proposes to use ordinal (not binary) correctness values to capture the ambiguity in the quality of a generation. In a similar flavor, Baan et al. (2022) studies the issues in the evaluation of calibration (against the human majority) when there is intrinsic human disagreement on the label.
Remarks. Existing literature sometimes uses the terms uncertainty and confidence interchangeably. They do often seemingly coincide: When a model’s prediction has low confidence, we naturally consider this as a high uncertainty case. This, however, is treating $U(\mathbf{x})=-\max_{\mathbf{s}}C(\mathbf{x},\mathbf{s})$ as an uncertainty estimate. In general, a model may exhibit high uncertainty over its output space but still express high confidence in a specific output. Conversely, a model could have low overall uncertainty but low confidence in a particular prediction. While the “low uncertainty low confidence case” is relatively less interesting in classification or regression tasks due to MLE point prediction, this scenario is notably more common in NLG, as the output is typically randomly sampled [^1] In fact, even if the output is greedily generated, it might not have the highest confidence as measured by Eq. 1. from the predictive distribution. There are also applications that require one but not the other (e.g. conformal language modeling (Quach et al., 2023) or seletive generation (Cole et al., 2023)). In the rest of this paper, we sometimes follow the language of the original papers and treat confidence estimates as uncertainty, but will clearly mark the methods that provide confidence estimates.
3. UQ Methods for Different Dimensions
@@
Method Uncertainty Stages Efficency Features Access to model Confidence
@@
Uncertainty Quantification for In-Context Learning (Ling et al., 2024) Input Uncertainty Multi Rounds Generations Black-box No
@@
TopologyUQ (Da et al., 2025a) Reasoning Uncertainty Multi Rounds Generations Black-box No
@@
Supervised uncertainty estimation(Liu et al., 2024c) Parameter + Prediction Uncertainty Fine-tuning White-box Yes
@@
BLoRA (Yang et al., 2023) Parameter Uncertainty Fine-tuning White-box Yes
@@
Shifting Attention to Relevance (SAR) (Duan et al., 2024) Prediction Uncertainty Single Round Generation White-box Yes
@@
Conformal Prediction (Quach et al., 2023; Kumar et al., 2023) Prediction Uncertainty Multi Rounds Generations White-box No
@@
Level-adaptive conformal prediction (Cherian et al., 2025) Prediction Uncertainty Multi Rounds Generations White-box No
@@
Kernel Semantic Entropy (Nikitin et al., 2024) Prediction Uncertainty Multi Rounds Generations + Additional Model White-box Yes
@@
MD-UQ (Chen et al., 2025) Prediction Uncertainty Multi Rounds Generations + Additional Model Black-box No
U-UE (Da et al., 2024a) Prediction Uncertainty Multi Rounds Generations + Additional Model Black-box Yes
Table 2. An overview of uncertainty quantification methods discussed in this paper.
@@ 3.1. Input Uncertainty
As mentioned in Section 2.1.2, input uncertainty arises from the ambiguous or incomplete input to the LLMs and many works in the LLMs domain try to deal with ambiguity. For example, there are many datasets on ambiguity with different tasks such as ambiguity detection (Min et al., 2020a; Guo et al., 2021) and many works on dealing with ambiguity (Zamani et al., 2020; Deng et al., 2023). However, these works did not consider uncertainty at all.
As far as we know, most uncertainty quantification methods that care about input uncertainty focus on perturbing the input prompts of LLMs. For instance, Hou et al. (2024) proposes an approach that generates multiple clarifications for a given prompt and ensembles the resulting generations by using mutual information to capture the disagreement among the predictions arising from different clarifications. Similarly, Ling et al. (2024) quantified the input uncertainty in the setting of in-context learning by using different in-context samples. Gao et al. (2024) proposes SPUQ, which perturbs the input by techniques such as paraphrasing and dummy tokens to expose the model’s sensitivity and capture uncertainty. Specifically, SPUQ quantified the input uncertainty by using a similarity metric such as BERTScore (Zhang et al., 2019) to measure how consistent the responses are across different perturbations.
We can see that there are only a few papers that care about input uncertainty and its application to ambiguity. Since ambiguity is common and important in natural language, we call more effort into input uncertainty and its application.
@@ 3.2. Reasoning Uncertainty
Recent research has advanced stepwise uncertainty quantification in LLM reasoning by explicitly eliciting and analyzing the internal reasoning process. For example, Tree of Uncertain Thoughts (TouT) (Mo and Xin, 2024) extends the Tree of Thoughts (ToT) (Yao et al., 2023) framework by addressing the inherent local uncertainties that occur during intermediate reasoning steps. TouT leverages Monte Carlo Dropout to assign uncertainty scores to pivotal decision points and, by integrating these local measures with global search techniques, it enhances the precision of response generation. Similarly, TopologyUQ (Da et al., 2025a) introduces a formal method to extract and structure LLM explanations into graph representations, quantifying reasoning uncertainty by employing graph-edit distances and revealing redundancy through stable topology measures. In addition, Uncertainty-aware Adaptive Guidance (UAG) (Yin et al., 2024) tackles error accumulation in multi-step reasoning by monitoring the predicted probability of the next token at each generation step, dynamically retracting to more reliable states and incorporating certified reasoning clues when high uncertainty is detected. Complementing these methods, Stable-Explanation Confidence (Becker and Soatto, 2024) quantifies uncertainty by examining the distribution of generated explanations, treating each model’s explanation pair as a test-time classifier to construct a posterior answer distribution that reflects overall reasoning confidence. More recently, CoT-UQ (Zhang and Zhang, 2025) has integrated chain-of-thought reasoning into a response-level uncertainty quantification framework, thereby leveraging the inherent multi-step reasoning capability of LLMs to further improve uncertainty assessment. Collectively, these approaches provide a robust and interpretable framework for enhancing LLM reasoning by quantifying uncertainty at both local and global levels.
@@ 3.3. Parameter Uncertainty
Parameter uncertainty arises when an LLM lacks sufficient knowledge due to limitations in its training data or model parameters. Parameter uncertainty reflects the model’s uncertainty about its own predictions, which can be reduced with additional training or better adaptation techniques.
Traditional UQ methods like Monte Carlo (MC) Dropout and Deep Ensembles have been widely used but are computationally infeasible for large-scale LLMs due to the need for multiple forward passes or model replicas. To address this, Bayesian Low-Rank Adaptation by Backpropagation (BLoB)(Wang et al., 2025) and Bayesian Low-Rank Adaptation (BLoRA)(Yang et al., 2023) incorporate Bayesian modeling into LoRA adapters, allowing uncertainty estimation through parameter distributions without full-model ensemble. However, these methods still incur significant computational costs.
@@ 3.4. Prediction Uncertainty
Most off-the-shelf uncertainty quantification methods focus on prediction uncertainty since it is the most straightforward way to estimate the uncertainty.
@@ 3.4.1. Single Round Generation
Most single-round generation methods utilize the logit or hidden states during the generation procedure. Single-round generation methods have the highest efficiency.
Perplexity is a measure of how well a probabilistic language model predicts a sequence of text (Vaswani et al., 2017) while Mora-Cross and Calderon-Ramirez (2024), Margatina et al. (2023) and Manakul et al. ([n. d.]) utilize the perplexity as the uncertainty. In detail, using $w_{i}$ as the i-th token in the generation, perplexity is given by:
(3) $\text{Perplexity}=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\ln p(w_{i})\right)$
A higher perplexity means the model spreads its probability more broadly over possible words, indicating that it has a higher uncertainty.
Maximum Token Log-Probability. Apart from the perplexity, Manakul et al. ([n. d.]) also uses maximum token log-probability:
(4) $Max(p)=\max\limits_{i}(-\ln p(w_{i}))$
Maximum token log-probability measures the sentence’s likelihood by assessing the least likely token in the sentence. A higher $\text{Maximum}(p)$ indicates higher uncertainty of the whole generation.
Entropy reflects how widely distributed a model’s predictions are for a given input, indicating the level of uncertainty in its outputs (Malinin and Gales, 2018; Kuhn et al., 2023; Kadavath et al., 2022).Entropy for the i-th token is provided by:
(5) $\mathcal{H}_{i}=-\sum_{\tilde{w}\in\mathcal{D}}p_{i}(\tilde{w})\log p_{i}(\tilde{w})$
Then it is possible to use the mean or maximum value of entropy as the final uncertainty (Manakul et al., [n. d.]):
(6) $Avg(\mathcal{H})=\frac{1}{N}\sum_{i=1}^{N}\mathcal{H}_{i};Max(\mathcal{H})=\max\limits_{i}(\mathcal{H}_{i})$
Furthermore, Shifting Attention to Relevance (SAR), proposed by Duan et al. (2024), enhanced the performance of entropy by adjusting attention to more relevant tokens inside the sentence. In detail, SAR assigned weight for $\mathcal{H}_{i}$ and the weight $R(w_{i},s,x)$ can be obtained by:
(7) $R(w_{i},s,x)=1-\left|g(x\cup s,x\cup s\setminus\{w_{i}\})\right|,$
Where $g$ is a function that measures the semantic similarity between two sentences such as NLI models.
Response Improbability. Fadeeva et al. (2024) uses response improbability, which computes the probability of a given sentence and subtracts the resulting value from one. In detail, response improbability is provided by:
(8) $MP(s)=1-\prod_{i=1}p_{i}(w_{i}).$
If the sentence is very certain (i.e., the product of token probabilities is high), $MP(s)$ will be low.
P(True) Kadavath et al. (2022) proposes P(True) [^2] The original name is P(IK), which stands for “I Know”., which measures the uncertainty of the claim by asking the LLM itself whether the generation is true or not. Specifically, P(True) is calculated:
@@
Note that here we are using $y_{1}$ as the first token instead of $w_{1}$ because $w_{1}$ represents the first token to the generation $s$ while $y_{1}$ represented the first token when asking LLM whether the generation $s$ is correct or not. P(True) requires to run the LLM twice. However, it does not require multiple generations $s$. Therefore, we still classify this method to single-round generation. (This could be considered an uncertainty estimate as the sequence to be evaluated is the prediction given the input.)

### v2

@@ 2.2.3. Confidence Estimation for LLMs
$\bullet$ UQ methods with Confidence Estimation. As uncertainty and confidence are often intertwined, many approaches used in UQ have their counterpart in confidence estimation. For example, for black-box settings where the parameters of LLMs are unavailable, (Lin et al., 2023; Zhang et al., 2024b) computes a similarity matrix of sampled responses and derives confidence estimates for each generation via its degree or distance derived from the graph Laplacian, before using these scores to compute uncertainty. For white-box settings where model parameters are available, researchers mostly compute the confidence from the output logits, either through normalizing Eq. 1 with the length of $\mathbf{s}$  (Malinin and Gales, 2021), replacing the logit-sum or mean with weighted sum by attention values (Lin et al., 2024) or by importance inferred from natural language inference (NLI) models (Duan et al., 2024). Such variants of sequence likelihood could then be fed for (entropy-style) uncertainty computation (Kuhn et al., 2023; Lin et al., 2024).
@@
$\bullet$ LLM-as-a-judge. Another popular approach is asking the LM itself whether a particular free-form generation is correct (Kadavath et al., 2022). However, this formulation also poses a restriction on the confidence estimation method, as it is essentially a scalar logit. Thus, many extensions focus on applying calibration methods from classification to calibrate such self-evaluation. The few exceptions include (Ren et al., 2023b; Kadavath et al., 2022), which converts samples from free-form generation into a multiple-choice question (with generations being the options) and adds a ”None of the above” option to elicit the confidence.
@@
$\bullet$ Trainable Confidence Estimators. Since we typically care about the LM’s confidence in the “semantic space” due to semantic invariance, instead of manipulating logits, a popular approach is to perform additional training for confidence estimation. This could be done on the base LM (either fully (Zhao et al., 2023a; Jiang et al., 2021; Kapoor et al., 2024) or partially (Liu et al., 2024a)) with a different loss, or using a separate model on the internal or external representations from the base LM (Jagannatha and Yu, 2020; Ulmer et al., 2024). On the other end of the spectrum, without any training, prompting could be used to elicit verbalized confidence values (Tian et al., 2023). Finally, one could combine multiple confidence estimation methods and enjoy the benefit of ensembling (Gao et al., 2024).
@@
As with UQ evaluation (more in Section 4), the choice of correctness function has a profound impact on the conclusion of the experiments, especially for free-form generation tasks. Popular choices include using (potentially larger) LLM as judges (Liu et al., 2024a; Lin et al., 2023; Tian et al., 2023), human annotations (Ren et al., 2023b), or lexical similarities such as ROUGE (Kuhn et al., 2023; Zhao et al., 2023a). Recently, Liu et al. (2025) proposes to evaluate free-form generation confidence measures with selected multiple-choice datasets as an efficient complement. For longer generations, Huang et al. (2024a) proposes to use ordinal (not binary) correctness values to capture the ambiguity in the quality of a generation. In a similar flavor, (Baan et al., 2022) studies the issues in the evaluation of calibration when there is intrinsic human disagreement on the label.
@@
Method Uncertainty Dimensions Efficency Features Access to Model Confidence
@@
ICL-Sample (Ling et al., 2024) Input Uncertainty Multi Rounds Generations Black-box No
@@
TopologyUQ (Da et al., 2025) Reasoning Uncertainty Multi Rounds Generations Black-box No
@@
Supervised estimation(Liu et al., 2024c) Parameter + Prediction Uncertainty Fine-tuning White-box Yes
@@
BLoRA (Yang et al., [n. d.]) Parameter Uncertainty Fine-tuning White-box Yes
@@
SAR (Duan et al., 2024) Prediction Uncertainty Single Round Generation White-box Yes
@@
Conformal Prediction (CP) (Quach et al., [n. d.]; Kumar et al., 2023) Prediction Uncertainty Multi Rounds Generations White-box No
@@
Level-adaptive CP (Cherian et al., 2025) Prediction Uncertainty Multi Rounds Generations White-box No
@@
Kernel Language Entropy (Nikitin et al., 2024) Prediction Uncertainty Multi Rounds Generations + Additional Model White-box Yes
@@
MD-UQ (Chen et al., 2025b) Prediction Uncertainty Multi Rounds Generations + Additional Model Black-box No
D-UE (Da et al., 2024a) Prediction Uncertainty Multi Rounds Generations + Additional Model Black-box Yes
Table 2. An overview of UQ methods discussed in this paper for different dimensions, efficiency, and model settings.
Remarks. Existing literature sometimes uses the terms uncertainty and confidence interchangeably. They do often seemingly coincide: When a model’s prediction has low confidence, we naturally consider this as a high uncertainty case. This, however, is treating $U(\mathbf{x})=-\max_{\mathbf{s}}C(\mathbf{x},\mathbf{s})$ as an uncertainty estimate. In general, a model may exhibit high uncertainty over its output space but still express high confidence in a specific output. Conversely, a model could have low overall uncertainty but low confidence in a particular prediction. While the “low uncertainty low confidence case” is relatively less interesting in classification or regression tasks due to MLE point prediction, this scenario is notably more common in NLG, as the output is typically randomly sampled [^1] In fact, even if the output is greedily generated, it might not have the highest confidence as measured by Eq. 1. from the predictive distribution. There are also applications that require one but not the other (e.g. conformal language modeling (Quach et al., [n. d.]) or seletive generation (Cole et al., 2023)). In the rest of this paper, we sometimes follow the language of the original papers and treat confidence estimates as uncertainty, but will clearly mark the methods that provide confidence estimates.
3. UQ Methods for Different Dimensions
@@ 3.1. Input Uncertainty
As mentioned in Section 2.1.2, input uncertainty arises from the ambiguous or incomplete input to the LLMs. While there are works in the LLMs domain that try to benchmark or deal with ambiguity (Guo et al., 2021; Zamani et al., 2020; Deng et al., 2023; Chen et al., 2025a), they did not model the uncertainties induced by ambiguity. Existing UQ methods that specifically consider input uncertainty focus on perturbing the input prompts of LLMs. For instance, (Hou et al., 2024) proposes an approach that generates multiple clarifications for a given prompt and ensembles the resulting generations by using mutual information to capture the disagreement among the predictions arising from different clarifications. Similarly, (Ling et al., 2024) proposed ICL-Sample, which quantified the input uncertainty in the setting of in-context learning using different in-context samples. (Gao et al., 2024) proposes SPUQ, which perturbs the input by techniques such as paraphrasing and dummy tokens to expose the model’s sensitivity and capture uncertainty. Specifically, SPUQ quantified the input uncertainty by using a similarity metric such as BERTScore (Zhang et al., [n. d.]) to measure how consistent the responses are across different perturbations. In general, there are only a few papers that consider input uncertainty. Since ambiguity is common and important in natural language, more effort is needed into input uncertainty and its application.
@@ 3.2. Reasoning Uncertainty
Reasoning is the process of drawing conclusions based on available information. As the LLMs have demonstrated remarkable performance on tasks involving reasoning, recent research has focused on using UQ in LLM reasoning and analyzing the internal reasoning process. For example, TopologyUQ (Da et al., 2025) introduces a formal method to extract and structure LLM explanations into graph representations, quantifying reasoning uncertainty by employing graph-edit distances and revealing redundancy through stable topology measures. Stable-Explanation Confidence (Becker and Soatto, 2024) treats each possible model and its explanation pair as a test-time classifier to construct a posterior answer distribution that reflects overall reasoning confidence. CoT-UQ (Zhang and Zhang, 2025) integrates chain-of-thought reasoning into a response-level UQ framework, thereby leveraging the inherent multi-step reasoning capability of LLMs to further improve uncertainty assessment. Collectively, these approaches provide a robust and interpretable framework for enhancing LLM reasoning by quantifying uncertainty at local or global levels.
The quantified uncertainty could be used to guide the exploration of reasoning steps and improving the final performance in completing the tasks. In (Mo and Xin, 2024), they propose Tree of Uncertain Thoughts (TouT), which extend the Tree of Thoughts (ToT) (Yao et al., 2023) framework by quantifying the uncertainties in intermediate reasoning steps with Monte Carlo Dropout and assigning uncertainty scores to important decision points. Similarly,  (Yin et al., 2024) reduces the error accumulation in multi-step reasoning by monitoring the predicted probability of the next token at each generation step, dynamically retracting to more reliable states and incorporating certified reasoning clues when high uncertainty is detected. Their experimental results shows that integrating uncertainty enhances the precision of generated responses by integrating these local measures with global search techniques.
@@ 3.3. Parameter Uncertainty
Parameter uncertainty arises when an LLM lacks sufficient knowledge due to limitations in its training data or model parameters. It reflects the model’s uncertainty about its own predictions, which can be reduced with additional training or adaptation techniques.
Traditional UQ methods like Monte Carlo Dropout and Deep Ensembles have been widely used but are computationally infeasible for large-scale LLMs due to the need for multiple forward passes or model replicas. To address this, Bayesian Low-Rank Adaptation by Backpropagation (BLoB) (Wang et al., 2025) and Bayesian Low-Rank Adaptation (BLoRA) (Yang et al., [n. d.]) incorporate Bayesian modeling into LoRA adapters, allowing uncertainty estimation through parameter distributions without a full-model ensemble. However, these methods still incur significant computational costs.
@@ 3.4. Prediction Uncertainty
Most off-the-shelf UQ methods focus on prediction uncertainty since it is the most straightforward way to estimate the uncertainty. Considering the number of generations and models when estimating uncertainties, existing methods for predicting uncertainty can be categorized into the following three categories.
@@ 3.4.1. Single Round Generation
Most single-round generation methods utilize the logit or hidden states during generation. With only one round of generation, these methods usually methods are usually efficient in estimating uncertainties.
$\bullet$ Perplexity is a measure of how well a probabilistic language model predicts a sequence of text (Vaswani et al., 2017) while Mora-Cross and Calderon-Ramirez (2024), Margatina et al. (2023) and Manakul et al. ([n. d.]) utilize the perplexity as the uncertainty. In detail, using $w_{i}$ as the i-th token in the generation, perplexity is given by $\text{Perplexity}=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\ln p(w_{i})\right)$. A higher perplexity means the model spreads its probability more broadly over possible words, indicating that it has a higher uncertainty.
$\bullet$ Maximum Token Log-Probability. Apart from the perplexity, Maximum token log-probability (Manakul et al., [n. d.]) measures the sentence’s likelihood by assessing the least likely token in the sentence. A higher $\text{Maximum}(p)$ indicates higher uncertainty of the whole generation. It is calculated by $Max(p)=\max\limits_{i}(-\ln p(w_{i}))$.
$\bullet$ Entropy reflects how widely distributed a model’s predictions are for a given input, indicating the level of uncertainty in its outputs (Kuhn et al., 2023; Kadavath et al., 2022). Entropy for the i-th token is provided by $\mathcal{H}_{i}=-\sum_{\tilde{w}\in\mathcal{D}}p_{i}(\tilde{w})\log p_{i}(\tilde{w})$. Then it is possible to use the mean or maximum value of entropy as the final uncertainty (Manakul et al., [n. d.]): $Avg(\mathcal{H})=\frac{1}{N}\sum_{i=1}^{N}\mathcal{H}_{i};Max(\mathcal{H})=\max\limits_{i}(\mathcal{H}_{i})$. Furthermore, Shifting Attention to Relevance (SAR) (Duan et al., 2024), enhanced the performance of entropy by adjusting attention to more relevant tokens inside the sentence. In detail, SAR assigned weight for $\mathcal{H}_{i}$ and the weight $R(w_{i},s,x)$ can be obtained by: $R(w_{i},s,x)=1-\left|g(x\cup s,x\cup s\setminus\{w_{i}\})\right|,$ where $g$ is a function that measures the semantic similarity between two sentences, which can be estimated with NLI models (Duan et al., 2024).
$\bullet$ Response Improbability (Fadeeva et al., 2024) uses response improbability, which computes the probability of a given sentence and subtracts the resulting value from one. In detail, response improbability is provided by $MP(s)=1-\prod_{i=1}p_{i}(w_{i}).$ If the sentence is certain (i.e., the product of token probabilities is high), $MP(s)$ will be low.